OpenStudy (anonymous):

A Kite 70 feet above the ground moves horizontally at a speed of 4 ft/s. At what rate is the angle between the string and the ground decreasing when 150 feet of string is out?

OpenStudy (johnweldon1993):

Well diagram it out first 

OpenStudy (johnweldon1993):

You know it is moving horizontally at 4ft/s so \(\large \frac{dx}{dt} = 4\)

OpenStudy (johnweldon1993):

Any thoughts from here? or still stuck?

OpenStudy (anonymous):

everything you just said is as far as I get

OpenStudy (johnweldon1993):

Okay, so what we want next, is some expression for 'x' in terms of theta Why? Well remember we eventually want to solve for \(\large \frac{d\theta}{dt}\) So hmm...I see a relation \(\large tan(\theta) = \frac{70}{x}\) right?

OpenStudy (johnweldon1993):

If we solve that for 'x'...we have \[\large x = \frac{70}{tan(\theta)}\] right?

OpenStudy (anonymous):

right

OpenStudy (johnweldon1993):

Okay good, so lets take the derivative of that wrt time Oh we cant just yet...we need to use the chain rule \[\large \frac{dx}{dt} = \frac{dx}{d\theta} \frac{d\theta}{dt}\] Now we can

OpenStudy (anonymous):

where did that equation come from?

OpenStudy (johnweldon1993):

That's just me rewriting something Notice \[\large \frac{dx}{dt} = \frac{dx}{\cancel{d\theta}}\frac{\cancel{d\theta}}{dt} = \frac{dx}{dt}\] Not changing anything...but at the SAME time...it is allowing us to execute our derivative

OpenStudy (anonymous):

oh ok

OpenStudy (johnweldon1993):

So now, what is the derivative with respect to theta? \[\large \frac{dx}{d\theta} \frac{70}{tan(\theta)} = ?\]

OpenStudy (johnweldon1993):

Sorry I actually gotta get going so I'm gonna do this out quick, I'll check back in a little bit to see how you did or if any questions arose So Recap: \[\large \frac{dx}{dt} = 4\] Now take derivative of \[\large x = \frac{70}{tan(\theta)}\] With respect to theta \[\large \frac{dx}{d\theta}\frac{70}{tan(\theta)} = -70csc^2(\theta)\] Now remember where I had that new way of writing the derivative? \[\large \frac{dx}{dt} = \frac{dx}{d\theta} \frac{d\theta}{dt}\] Plug in what we have \[\large 4 = -70csc^2(\theta)\frac{d\theta}{dt}\] Here, it is important to see at this moment we are solving for, the opposite of theta is 70 and the hypotenuse is 150...why is that important? Well \[\large 4 = -70(\frac{1}{sin^2(\theta)})\frac{d\theta}{dt} \] We know that at that exact moment...sin(theta) = 70/150 = 1/2 So that means sin^2(theta) = 1/4 So what we have now is \[\large 4 = -70(\frac{1}{\frac{1}{4}})\frac{d\theta}{dt}\] Which yes is \[\large 4 = -70(4)\frac{d\theta}{dt}\] Meaning \[\large \frac{d\theta}{dt} = -\frac{4}{280} = -\frac{1}{70}\] Which is some decimal which is your answer *I'll check back in in a bit to see what you have :)

OpenStudy (anonymous):

it wasn't right